# Feynman Formula for Discrete-time Quantum Walks

### v1

@@ 5. Extensions to other quantum walks
In this section, we propose generalization of the Feynman formula (Theorem 2.7) by considering one-dimensional quantum walks with other coin operators, which include more general homogeneous coins, time-dependent coins (see, for example, [Bañuls et al., 2006]) and site-dependent coins (see, for example, [Wójcik et al., 2004], [Maeda and Suzuki, 2020], [Maeda et al., 2018] and [Linden and Sharam, 2009]), and discuss, if possible, their asymptotic behavior and limiting PDEs. We show that all such quantum walks admit a Feynman-type formula, which justifies the robustness of writing a Feynman-type formula.
@@ 5.1. General unitary coins
(5.1) $e^{i\phi}Z_{\beta}e^{i\theta\sigma_{1}}Z_{\alpha}=e^{i\phi}\begin{bmatrix}\beta&0\\ 0&\beta^{-1}\end{bmatrix}\begin{bmatrix}\cos\theta&i\sin\theta\\ i\sin\theta&\cos\theta\end{bmatrix}\begin{bmatrix}\alpha&0\\ 0&\alpha^{-1}\end{bmatrix},\quad\mbox{with}\quad Z_{\omega}:=\begin{bmatrix}\omega&0\\ 0&\omega^{-1}\end{bmatrix}\quad\forall\omega\in\mathbb{C},$
@@
(5.3) $\psi_{n}^{\phi;\beta,\alpha;\theta}(x,z)=e^{in\phi}\sum_{k_{1},\cdots,k_{n}=0}^{\infty}$ $\frac{(i\theta)^{k_{1}+\cdots+k_{n}}}{k_{1}!\cdots k_{n}!}\alpha^{(-1)^{k_{1}}z+\cdots+(-1)^{k_{1}+\cdots+k_{n}}z}\beta^{z+(-1)^{k_{1}}z+\cdots+(-1)^{k_{1}+\cdots+k_{n-1}}}$
$\times\psi_{0}(x-z-(-1)^{k_{1}}z-\cdots-(-1)^{k_{1}+\cdots+k_{n-1}}z,z(-1)^{k_{1}+\cdots+k_{n}}z),$
@@
(5.11) $\psi(t,y,z,\eta,B,A,\lambda):=\lim_{\epsilon\to 0^{+}}\psi_{[t/\epsilon]}^{\epsilon\eta;e^{i\epsilon B},e^{i\epsilon A};\epsilon\lambda,\epsilon}([y/\epsilon],z)=e^{it\eta+t\lambda}\mathbb{E}\left[i^{N_{t}}e^{i(A+B)zT_{t}}f(y-zT_{t},z(-1)^{N_{t}})\right],$
@@
(5.13) $\psi(t+h,y,z,\eta,B,A,\lambda)=e^{h\lambda+ih\eta}\mathbb{E}\left[i^{N_{h}}e^{i(A+B)zT_{h}}\psi(t,y-zT_{h},z(-1)^{N_{h}},\eta,B,A,\lambda)\right],$
@@ 5.2.1. Rotation coin with site-dependent global phase
(5.17) $\psi_{n}^{V;\theta}(x,z)=\sum_{k_{1},\cdots,k_{n}=0}^{\infty}$ $\frac{(i\theta)^{k_{1}+\cdots+k_{n}}}{k_{1}!\cdots k_{n}!}e^{i[V(x-z)+V(x-z-(-1)^{k_{1}}z)-\cdots-V(x-z-(-1)^{k_{1}}z-\cdots-(-1)^{k_{n-1}}z)]}$
@@
(5.18) $\psi(t,y,z):=\lim_{\epsilon\to 0}\psi_{[t/\epsilon]}^{\epsilon V(\cdot\epsilon),\epsilon\lambda;\epsilon}([y/\epsilon],z)=e^{t\lambda}\mathbb{E}\left[i^{N_{t}}e^{i\int_{0}^{t}V(y-zT_{s})ds)}f(y-zT_{t},z(-1)^{N_{t}})\right],$
@@ 5.3. Time-dependent coins
(5.25) $\psi_{n+1}=W_{n+1}\psi_{n},\quad\psi_{0}\in\ell_{2}(\mathbb{Z})\otimes\ell_{2}(\pm 1),$
@@ 5.4. Yet another walk
The previous result has a continuous-time version which was known for some time as Molchanov formula [Carmona, 1984]. Let $(X_{t})_{t\geq 0}$ be the continuous-time Markov chain with transition matrix $M$ and $(N_{t})_{t\geq 0}$ be the Poisson process with intensity $\lambda>0$ describing the jump times of the chain. Given a bounded potential $V(x)_{x\in\mathbb{Z}}$, we have the following:
@@ 6.2. Outlook
In Section 5, we demonstrated the robustness of the Feynman formula on inhomogeneous coins; yet these coins can yield non-ballistic behaviors, including localized and hybrid behaviors, as seen in the numerical experiments by [Ahmad et al., 2020] for site-dependent coins and [Panahiyan and Fritzsche, 2018] for time-dependent coins. It is natural to ask how these properties are reflected in the associated Markov additive process. Meanwhile, quantum walks with higher-dimensional state space and chirality space allow one to study entanglement effects involved in a quantum walk (see [Watabe et al., 2008] which discussed quantum walks on $\mathbb{Z}^{2}$ and [Venegas-Andraca et al., 2005] which discussed one-dimensional quantum walks with higher dimensional, entangled coins). It is natural to extend the robustness of the Feynman formula to these higher-dimensional generalizations. Lastly, we are also currently working on probabilistic representations for the nonlinear quantum walks and their associated Dirac PDEs described in [Maeda et al., 2018] and [Maeda and Suzuki, 2020].
Appendix A Sketch of proof of Lemma 3.6
The method of stationary phase (see for instance [Olver, 1997] and [Wong, 2001] for details) entails that Fourier integrals of the form
@@
where $\{k_{j}\}_{j=1}^{N}$ are critical points to $\Theta(k)$ that are interior to $[a,b]$ for some $N\in\mathbb{N}$ under certain regularity assumption on $g$ and $\Theta$. In this section we would use the method of stationary phase to both derive (3.20), (3.21) and (3.22) formally and to justify all assumptions we made in the formal computation in A.1.3.
A.1. The limit (3.20)
A.1.1. Formal computation of single application of stationary phase
We start with by formally applying the method of stationary phase for a fix $a\in(-\cos\theta+\epsilon,\cos\theta-\epsilon)$ on (3.20), assuming that we can replace $\left\lceil an\right\rceil$ with $an$, when $n$ is large enough (see A.1.3 for a justification):
@@ A.1.1. Formal computation of single application of stationary phase
(A.11) $r_{B_{1}}(k,0)\ell_{B_{1}}(k,0)=\frac{1}{4}\Big(1+\frac{\sin k\cos\theta}{\sqrt{1-\cos^{2}\theta\cos^{2}k}}\Big),$ $r_{B_{2}}(k,0)\ell_{B_{2}}(k,0)=\frac{1}{4}\Big(1-\frac{\sin k\cos\theta}{\sqrt{1-\cos^{2}\theta\cos^{2}k}}\Big),$
@@
(A.12) $\frac{1}{2\pi}\frac{1+a}{4}\sqrt{\frac{2\pi^{2}}{n}f_{K}(a;\theta)}\times$ $\{e^{in\cos^{-1}(\cos k_{1}\cos\theta)-ik_{1}an+i\operatorname{sgn}(a)\frac{\pi}{4}}+e^{in\cos^{-1}(\cos k_{2}\cos\theta)-ik_{2}an-i\operatorname{sgn}(a)\frac{\pi}{4}}$
@@
(A.13) $2\cos(n\Theta_{B_{1}}(k_{1},a)+\operatorname{sgn}(a)\frac{\pi}{4})+2\cos(n\Theta_{B_{1}}(k_{2},a)-\operatorname{sgn}(a)\frac{\pi}{4})$
$=$ $4\cos(\frac{n}{2}[\Theta_{B_{1}}(k_{1},a)+\Theta_{B_{1}}(k_{2},a)])\cos(\frac{n}{2}[\Theta_{B_{1}}(k_{1},a)-\Theta_{B_{2}}(k_{2},a)]+\operatorname{sgn}(a)\frac{\pi}{4})$
@@
$=$ $\frac{4}{\sqrt{2}}\cos(\frac{n\pi(1-a)}{2})\Big[\cos(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})-\operatorname{sgn}(a)\sin(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})\Big],$
@@
(A.15) $\frac{1+a}{2\sqrt{n}}\sqrt{f_{K}(a;\theta)}\cos(\frac{n\pi(1-a)}{2})\Big[\cos(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})-\operatorname{sgn}(a)\sin(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})\Big]$
A.1.2. Formal computation for integrating leading terms
In the previous section, we formally showed that $e^{n\theta}\sum_{j=1,2}I_{n}(0,an,B_{j})$ is asymptotically equivalent to (A.15); in this section we continue from there to formally verify 3.20. To begin with, taking modulus square on(A.15) gives the asymptotic relation:
@@ A.1.2. Formal computation for integrating leading terms
Note that this asymptotic relation is valid because the domain of integral is contained in $(-\cos\theta+\epsilon,\cos\theta-\epsilon)$ which is away from the problematic boundary points $\pm\cos\theta$. Due to the existence of fast oscillation terms, as $n\to\infty$, the last three integrals in the asymptotic equivalence converge to 0 by both the Riemann-Lebesgue Lemma and the method of stationary phase (see A.1.3 for details). The first term converges and we have
@@ A.1.3. Rigorous justification
Because of the assumption that $a\in(-\cos\theta+\epsilon,\cos\theta-\epsilon)$, we can in fact have the large $n$ asymptotic relation (A.5) with $a$ replaced by $a_{n}:=\left\lceil an\right\rceil/n$, which is uniform across all choices of $a$. The precise argument has been made in Proposition 2.2 of [Sunada and Tate, 2012]. Nonetheless, the formal procedure of using $a$ and $a_{n}$ yield similar computations.
@@ A.2. The remaining two limits
(A.29) $r_{B_{1}}(k,0)\ell_{B_{1}}(k,0)=\frac{1}{4i}\frac{\cos k\sin\theta+i\sin k\sin\theta}{\sqrt{1-\cos^{2}k\cos^{2}\theta}},$ $r_{B_{2}}(k,0)\ell_{B_{2}}(k,0)=-\frac{1}{4i}\frac{\cos k\sin\theta+i\sin k\sin\theta}{\sqrt{1-\cos^{2}k\cos^{2}\theta}}.$
@@
(A.30) $\qquad\frac{\cos k_{1}\sin\theta}{\sqrt{1-\cos^{2}k_{1}\cos^{2}\theta}}=\operatorname{sgn}(a)\sec\theta\sqrt{\cos^{2}\theta-a^{2}},$ $\frac{\sin k_{1}\sin\theta}{\sqrt{1-\cos^{2}k_{1}\cos^{2}\theta}}=a\tan\theta.$
Using (A.9) and (A.29) with the corresponding values at the saddle points given by (A.10) and (A.30) shows that $e^{n\theta}\sum_{j=1,2}I_{n}(1,an,B_{j})$ is asymptotically equivalent to:
(A.31) $\frac{1}{2\pi}\frac{\operatorname{sgn}(a)\sqrt{\cos^{2}\theta-a^{2}}\sec\theta}{4i}\sqrt{\frac{2\pi^{2}}{n}f_{K}(a;\theta)}$
@@
(A.32) $\frac{1}{2\pi}\frac{a\tan\theta}{4}\sqrt{\frac{2\pi^{2}}{n}f_{K}(a;\theta)}$
@@
$+e^{-in\cos^{-1}(\cos k_{3}\cos\theta)-ik_{3}an+i\operatorname{sgn}(a)\frac{\pi}{4}}+e^{-in\cos^{-1}(\cos k_{4}\cos\theta)-ik_{4}an-i\operatorname{sgn}(a)\frac{\pi}{4}}\},$
which is a sum of the two terms, (A.31) and (A.32). The oscillation term in (A.31) is given by
(A.33) $2\cos(n\Theta_{B_{1}}(k_{1},a)+\operatorname{sgn}(a)\frac{\pi}{4})-2\cos(n\Theta_{B_{1}}(k_{2},a)-\operatorname{sgn}(a)\frac{\pi}{4})$
$=$ $-4\sin(\frac{n\pi(1-a)}{2})\sin(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2}+\operatorname{sgn}(a)\frac{\pi}{4})$
$=$ $-\frac{4}{\sqrt{2}}\sin(\frac{n\pi(1-a)}{2})[\sin(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})+\operatorname{sgn}(a)\cos(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})]$
@@
$2\cos(n\Theta_{B_{1}}(k_{1},a)+\operatorname{sgn}(a)\frac{\pi}{4})+2\cos(n\Theta_{B_{1}}(k_{2},a)-\operatorname{sgn}(a)\frac{\pi}{4})$
$=$ $\frac{4}{\sqrt{2}}\cos(\frac{n\pi(1-a)}{2})\Big[\cos(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})-\operatorname{sgn}(a)\sin(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})\Big].$
@@
(A.34) $i\frac{\operatorname{sgn}(a)\sqrt{\cos^{2}\theta-a^{2}}\sec\theta}{2\sqrt{n}}\sqrt{f_{K}(a;\theta)}$
$\times\sin(\frac{n\pi(1-a)}{2})\Big[\sin(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})+\operatorname{sgn}(a)\cos(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})\Big]$
$+\frac{a\tan\theta}{2\sqrt{n}}\sqrt{f_{K}(a;\theta)}$
$\times\cos(\frac{n\pi(1-a)}{2})\Big[\cos(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})-\operatorname{sgn}(a)\sin(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})\Big].$
Taking modulus square on (A.34) gives the asymptotic equivalence of $|e^{2n\theta}\absolutevalue{\sum_{j=1,2}I_{n}(1,an,B_{j})}$:
@@
$=\int_{y_{1}}^{y_{2}}\frac{1}{8}f_{K}(a;\theta)((\cos^{2}\theta-a^{2})\sec^{2}\theta+a^{2}\tan^{2}\theta)\mathrm{d}a$
$=\frac{1}{8}\int_{y_{1}}^{y_{2}}(1-a^{2})f_{K}(a;\theta)\mathrm{d}a.$
A.2.2. Formal computation for (3.22)
Using the asymptotic equivalence obtained in (A.15) and (A.34), we deduce the asymptotic equivalence of $e^{2n\theta}(\sum_{j=1,2}I_{n}(0,an,B_{j}))\overline{(\sum_{j=1,2}I_{n}(1,an,B_{j}))}$, which is given by:

### v2

@@ 5. Extensions to other quantum walks
In this section, we propose generalization of the Feynman formula (Theorem 2.7) by considering one-dimensional quantum walks with other coin operators, which include more general homogeneous coins, time-dependent coins (see, for example, Bañuls et al., [2006]) and site-dependent coins (see, for example, Wójcik et al., [2004], Maeda and Suzuki, [2020], Maeda et al., [2018] and Linden and Sharam, [2009]), and discuss, if possible, their asymptotic behavior and limiting PDEs. We show that all such quantum walks admit a Feynman-type formula, which justifies the robustness of writing a Feynman-type formula.
@@ 5.1. General unitary coins
(5.1) $e^{i\phi}Z_{\beta}e^{i\theta\sigma_{1}}Z_{\alpha}$ $=e^{i\phi}\begin{bmatrix}\beta&0\\ 0&\beta^{-1}\end{bmatrix}\begin{bmatrix}\cos\theta&i\sin\theta\\ i\sin\theta&\cos\theta\end{bmatrix}\begin{bmatrix}\alpha&0\\ 0&\alpha^{-1}\end{bmatrix}$,
(5.2) $\quad\mbox{with}\quad Z_{\omega}$ $:=\begin{bmatrix}\omega&0\\ 0&\omega^{-1}\end{bmatrix}\quad\forall\omega\in\mathbb{C},$
@@
(5.4) $\begin{split}&\psi_{n}^{\phi;\beta,\alpha;\theta}(x,z)\\ &=e^{in\phi}\sum_{k_{1},\cdots,k_{n}=0}^{\infty}\frac{(i\theta)^{k_{1}+\cdots+k_{n}}}{k_{1}!\cdots k_{n}!}\alpha^{(-1)^{k_{1}}z+\cdots+(-1)^{k_{1}+\cdots+k_{n}}z}\\ &\qquad\qquad\qquad\times\beta^{z+(-1)^{k_{1}}z+\cdots+(-1)^{k_{1}+\cdots+k_{n-1}}}\\ &\qquad\qquad\qquad\times\psi_{0}(x-z-(-1)^{k_{1}}z-\cdots-(-1)^{k_{1}+\cdots+k_{n-1}}z,z(-1)^{k_{1}+\cdots+k_{n}}z),\end{split}$
@@
(5.12) $\psi(t,y,z,\eta,B,A,\lambda)$ $:=\lim_{\epsilon\to 0^{+}}\psi_{[t/\epsilon]}^{\epsilon\eta;e^{i\epsilon B},e^{i\epsilon A};\epsilon\lambda,\epsilon}([y/\epsilon],z)$
(5.13) $\,=e^{it\eta+t\lambda}\mathbb{E}\left[i^{N_{t}}e^{i(A+B)zT_{t}}f(y-zT_{t},z(-1)^{N_{t}})\right],$
@@
(5.15) $\psi(t+h,y,z,\eta,B,A,\lambda)$
(5.16) $=$ $e^{h\lambda+ih\eta}\mathbb{E}\left[i^{N_{h}}e^{i(A+B)zT_{h}}\psi(t,y-zT_{h},z(-1)^{N_{h}},\eta,B,A,\lambda)\right],$
@@ 5.2.1. Rotation coin with site-dependent global phase
(5.20) $\psi_{n}^{V;\theta}(x,z)=\sum_{k_{1},\cdots,k_{n}=0}^{\infty}$ $\frac{(i\theta)^{k_{1}+\cdots+k_{n}}}{k_{1}!\cdots k_{n}!}$
(5.21) $\times e^{i[V(x-z)+V(x-z-(-1)^{k_{1}}z)-\cdots-V(x-z-(-1)^{k_{1}}z-\cdots-(-1)^{k_{n-1}}z)]}$
@@
(5.22) $\psi(t,y,z)$ $:=\lim_{\epsilon\to 0}\psi_{[t/\epsilon]}^{\epsilon V(\cdot\epsilon),\epsilon\lambda;\epsilon}([y/\epsilon],z)$
(5.23) $=e^{t\lambda}\mathbb{E}\left[i^{N_{t}}e^{i\int_{0}^{t}V(y-zT_{s})\mathrm{d}s}f(y-zT_{t},z(-1)^{N_{t}})\right],$
@@ 5.3. Time-dependent coins
(5.30) $\psi_{n}=W_{n}\psi_{n-1}\quad\mbox{for}\quad n\geq 1;\quad\psi_{0}\in\ell_{2}(\mathbb{Z})\otimes\ell_{2}(\pm 1),$
@@ 5.4. Yet another walk
The previous result has a continuous-time version which was known for some time as Molchanov formula (see Carmona, [1986]). Let $(X_{t})_{t\geq 0}$ be the continuous-time Markov chain with transition matrix $M$ and $(N_{t})_{t\geq 0}$ be the Poisson process with intensity $\lambda>0$ describing the jump times of the chain. Given a bounded potential $V(x)_{x\in\mathbb{Z}}$, we have the following:
@@ 6.2. Outlook
In Section 5, we demonstrated the robustness of the Feynman formula on inhomogeneous coins; yet these coins can yield non-ballistic behaviors, including localized and hybrid behaviors, as seen in the numerical experiments by Ahmad et al., [2020] for site-dependent coins and Panahiyan and Fritzsche, [2018] for time-dependent coins. It is natural to ask how these properties are reflected in the associated Markov additive process. Meanwhile, quantum walks with higher-dimensional state space and chirality space allow one to study entanglement effects involved in a quantum walk (see Watabe et al., [2008] which discussed quantum walks on $\mathbb{Z}^{2}$ and Venegas-Andraca et al., [2005] which discussed one-dimensional quantum walks with higher dimensional, entangled coins). It is natural to extend the robustness of the Feynman formula to these higher-dimensional generalizations. Lastly, we are also currently working on probabilistic representations for the nonlinear quantum walks and their associated Dirac PDEs described in Maeda et al., [2018] and Maeda and Suzuki, [2020].
Sketch of proof of Lemma 3.6
The method of stationary phase (see for instance Olver, [1997] and Wong, [2001] for details) entails that Fourier integrals of the form
@@
where $\{k_{j}\}_{j=1}^{N}$ are critical points to $\Theta(k)$ that are interior to $[a,b]$ for some $N\in\mathbb{N}$ under certain regularity assumption on $g$ and $\Theta$. In this section we would use the method of stationary phase to both derive (3.20), (3.21) and (3.22) formally and to justify all assumptions we made in the formal computation in A.3.3.
A.3. The limit (3.20)
A.3.1. Formal computation of single application of stationary phase
We start with by formally applying the method of stationary phase for a fix $a\in(-\cos\theta+\epsilon,\cos\theta-\epsilon)$ on (3.20), assuming that we can replace $\left\lceil an\right\rceil$ with $an$, when $n$ is large enough (see A.3.3 for a justification):
@@ A.3.1. Formal computation of single application of stationary phase
(A.11) $\begin{split}r_{B_{1}}(k,0)\ell_{B_{1}}(k,0)&=\frac{1}{4}\Big(1+\frac{\sin k\cos\theta}{\sqrt{1-\cos^{2}\theta\cos^{2}k}}\Big),\\ r_{B_{2}}(k,0)\ell_{B_{2}}(k,0)&=\frac{1}{4}\Big(1-\frac{\sin k\cos\theta}{\sqrt{1-\cos^{2}\theta\cos^{2}k}}\Big),\end{split}$
@@
(A.12) $\frac{1}{2\pi}\frac{1+a}{4}\sqrt{\frac{2\pi^{2}}{n}f_{K}(a;\theta)}$
(A.13) $\times$ $\{e^{in\cos^{-1}(\cos k_{1}\cos\theta)-ik_{1}an+i\operatorname{sgn}(a)\frac{\pi}{4}}+e^{in\cos^{-1}(\cos k_{2}\cos\theta)-ik_{2}an-i\operatorname{sgn}(a)\frac{\pi}{4}}$
(A.14) $+e^{-in\cos^{-1}(\cos k_{3}\cos\theta)-ik_{3}an+i\operatorname{sgn}(a)\frac{\pi}{4}}+e^{-in\cos^{-1}(\cos k_{4}\cos\theta)-ik_{4}an-i\operatorname{sgn}(a)\frac{\pi}{4}}\}.$
@@
(A.15) $\qquad 2\cos(n\Theta_{B_{1}}(k_{1},a)+\operatorname{sgn}(a)\frac{\pi}{4})+2\cos(n\Theta_{B_{1}}(k_{2},a)-\operatorname{sgn}(a)\frac{\pi}{4})$
(A.16) $=$ $4\cos(\frac{n}{2}[\Theta_{B_{1}}(k_{1},a)+\Theta_{B_{1}}(k_{2},a)])\cos(\frac{n}{2}[\Theta_{B_{1}}(k_{1},a)-\Theta_{B_{2}}(k_{2},a)]+\operatorname{sgn}(a)\frac{\pi}{4})$
@@
$=$ $\frac{4}{\sqrt{2}}\cos(\frac{n\pi(1-a)}{2})\Big[\cos(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})$
$\qquad\qquad\qquad\qquad\qquad\qquad{}-\operatorname{sgn}(a)\sin(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})\Big],$
@@
(A.18) $\begin{split}&\frac{1+a}{2\sqrt{n}}\sqrt{f_{K}(a;\theta)}\cos(\frac{n\pi(1-a)}{2})\\ \times&\Big[\cos(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})-\operatorname{sgn}(a)\sin(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})\Big].\end{split}$
A.3.2. Formal computation for integrating leading terms
In the previous section, we formally showed that $e^{n\theta}\sum_{j=1,2}I_{n}(0,an,B_{j})$ is asymptotically equivalent to (A.18); in this section we continue from there to formally verify (3.20). To begin with, taking modulus square on (A.18) gives the asymptotic relation:
@@ A.3.2. Formal computation for integrating leading terms
Note that this asymptotic relation is valid because the domain of integral is contained in $(-\cos\theta+\epsilon,\cos\theta-\epsilon)$ which is away from the problematic boundary points $\pm\cos\theta$. Due to the existence of fast oscillation terms, as $n\to\infty$, the last three integrals in the asymptotic equivalence converge to 0 by both the Riemann-Lebesgue Lemma and the method of stationary phase (see A.3.3 for details). The first term converges and we have
@@ A.3.3. Rigorous justification
Because of the assumption that $a\in(-\cos\theta+\epsilon,\cos\theta-\epsilon)$, we can in fact have the large $n$ asymptotic relation (A.5) with $a$ replaced by $a_{n}:=\left\lceil an\right\rceil/n$, which is uniform across all choices of $a$. The precise argument has been made in Proposition 2.2 of Sunada and Tate, [2012]. Nonetheless, the formal procedure of using $a$ and $a_{n}$ yield similar computations.
@@ A.4. The remaining two limits
(A.32) $\begin{split}r_{B_{1}}(k,0)\ell_{B_{1}}(k,0)&=\frac{1}{4i}\frac{\cos k\sin\theta+i\sin k\sin\theta}{\sqrt{1-\cos^{2}k\cos^{2}\theta}},\\ r_{B_{2}}(k,0)\ell_{B_{2}}(k,0)&=-\frac{1}{4i}\frac{\cos k\sin\theta+i\sin k\sin\theta}{\sqrt{1-\cos^{2}k\cos^{2}\theta}}.\end{split}$
@@
(A.33) $\begin{split}\frac{\cos k_{1}\sin\theta}{\sqrt{1-\cos^{2}k_{1}\cos^{2}\theta}}&=\operatorname{sgn}(a)\sec\theta\sqrt{\cos^{2}\theta-a^{2}},\\ \frac{\sin k_{1}\sin\theta}{\sqrt{1-\cos^{2}k_{1}\cos^{2}\theta}}&=a\tan\theta.\end{split}$
Using (A.9) and (A.32) with the corresponding values at the saddle points given by (A.10) and (A.33) shows that $e^{n\theta}\sum_{j=1,2}I_{n}(1,an,B_{j})$ is asymptotically equivalent to the sum:
(A.34) $\qquad\frac{1}{2\pi}\frac{\operatorname{sgn}(a)\sqrt{\cos^{2}\theta-a^{2}}\sec\theta}{4i}\sqrt{\frac{2\pi^{2}}{n}f_{K}(a;\theta)}$
@@
(A.35) $\qquad\frac{1}{2\pi}\frac{a\tan\theta}{4}\sqrt{\frac{2\pi^{2}}{n}f_{K}(a;\theta)}$
@@
The oscillation term in (A.34) is given by
$\begin{split}&2\cos(n\Theta_{B_{1}}(k_{1},a)+\operatorname{sgn}(a)\frac{\pi}{4})-2\cos(n\Theta_{B_{1}}(k_{2},a)-\operatorname{sgn}(a)\frac{\pi}{4})\\ =&-4\sin(\frac{n\pi(1-a)}{2})\sin(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2}+\operatorname{sgn}(a)\frac{\pi}{4})\\ =&{}-\frac{4}{\sqrt{2}}\sin(\frac{n\pi(1-a)}{2})\Big[\sin(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})\\ &{}+\operatorname{sgn}(a)\cos(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})\Big],\end{split}$
@@
$\begin{split}&2\cos(n\Theta_{B_{1}}(k_{1},a)+\operatorname{sgn}(a)\frac{\pi}{4})+2\cos(n\Theta_{B_{1}}(k_{2},a)-\operatorname{sgn}(a)\frac{\pi}{4})\\ =&4\cos(\frac{n\pi(1-a)}{2})\cos(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2}+\operatorname{sgn}(a)\frac{\pi}{4})\\ =&\frac{4}{\sqrt{2}}\cos(\frac{n\pi(1-a)}{2})\Big[\cos(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})\\ &\qquad\qquad\qquad\qquad\qquad\qquad{}-\operatorname{sgn}(a)\sin(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})\Big].\end{split}$
@@
(A.36) $\begin{split}\quad&i\frac{\operatorname{sgn}(a)\sqrt{\cos^{2}\theta-a^{2}}\sec\theta}{2\sqrt{n}}\sqrt{f_{K}(a;\theta)}\sin(\frac{n\pi(1-a)}{2})\\ \quad&\quad\times\Big[\sin(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})+\operatorname{sgn}(a)\cos(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})\Big]\\ \quad&+\frac{a\tan\theta}{2\sqrt{n}}\sqrt{f_{K}(a;\theta)}\cos(\frac{n\pi(1-a)}{2})\\ \quad&\quad\times\Big[\cos(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})-\operatorname{sgn}(a)\sin(n\Theta_{B_{1}}(k_{1},a)-\frac{n\pi(1-a)}{2})\Big].\end{split}$
Taking modulus square on (LABEL:eqns:_full_single_stationary_phase_case1) gives the asymptotic equivalence of $e^{2n\theta}\lvert{\sum_{j=1,2}I_{n}(1,an,B_{j})}\rvert$:
@@
$=\int_{y_{1}}^{y_{2}}\frac{1}{8}f_{K}(a;\theta)((\cos^{2}\theta-a^{2})\sec^{2}\theta+a^{2}\tan^{2}\theta)\mathrm{d}a=\frac{1}{8}\int_{y_{1}}^{y_{2}}(1-a^{2})f_{K}(a;\theta)\mathrm{d}a.$
A.4.2. Formal computation for (3.22)
Using the asymptotic equivalence obtained in the two equations (A.18) and (LABEL:eqns:_full_single_stationary_phase_case1), we can now deduce the asymptotic equivalence of the product $e^{2n\theta}(\sum_{j=1,2}I_{n}(0,an,B_{j}))\overline{(\sum_{j=1,2}I_{n}(1,an,B_{j}))}$ to be: